OpenStudy (anonymous):

Which equation represents the line that passes through the points (−3, 7) and (3, 3)?

OpenStudy (anonymous):

First find the slope m by using formula m = y2 -y1 / x2 -x1 then use that slope m & any one point from given points & use slope point form y - y1 = m (x - x1) to find equation ..